OpenStudy (anonymous):

The amplitude of F (x) = (1/2)sin(x/2) is 1/2 2 4 I would assume it is 1/2 since that is the |A|, but since there is (x/2) at the end, would that make it 2?

OpenStudy (anonymous):

But I wouldn't think so since (1/2)*(2/1) is 1. Right?

jimthompson5910 (jim_thompson5910):

In general, the amplitude of y = A*sin(x) is |A|

jimthompson5910 (jim_thompson5910):

In this case, A = 1/2 So |A| = |1/2| = 1/2

jimthompson5910 (jim_thompson5910):

So your first answer is correct

OpenStudy (anonymous):

Oh okay, thanks :) I appreciate it!

jimthompson5910 (jim_thompson5910):

np